### v2

# CausalMed: Causality-Based Personalized Medication Recommendation Centered on Patient health state

Abstract.
Medication recommendation systems are developed to recommend suitable medications tailored to specific patient. Previous researches primarily focus on learning medication representations, which have yielded notable advances. However, these methods are limited to capturing personalized patient representations due to the following primary limitations: (i) unable to capture the differences in the impact of diseases/procedures on patients across various patient health states; (ii) fail to model the direct causal relationships between medications and specific health state of patients, resulting in an inability to determine which specific disease each medication is treating. To address these limitations, we propose CausalMed, a patient health state-centric model capable of enhancing the personalization of patient representations. Specifically, CausalMed first captures the causal relationship between diseases/procedures and medications through causal discovery and evaluates their causal effects. Building upon this, CausalMed focuses on analyzing the health state of patients, capturing the dynamic differences of diseases/procedures in different health states of patients, and transforming diseases/procedures into medications on direct causal relationships. Ultimately, CausalMed integrates information from longitudinal visits to recommend medication combinations. Extensive experiments on real-world datasets show that our method learns more personalized patient representation and outperforms state-of-the-art models in accuracy and safety.
Medication Recommendation, Causal Inference, Electronic Health Record
submissionid: 137ccs: Information systems Information retrievalccs: Computing methodologies Causal reasoning and diagnosticsccs: Applied computing Health informatics
1. Introduction
Given the noticeable imbalance in healthcare resources’ supply and demand environment, there has been a growing emphasis on AI-based medical systems (Rosaci, 2007; Choi et al., 2016a; Bansal et al., 2014; Ma et al., 2023). Medication recommendation, as one of the domains, aims to provide personalized medication treatment plans for specific patients using data mining (Ma et al., 2018; Symeonidis et al., 2021; Wang et al., 2021) and deep learning (Shang et al., 2018; Wang et al., 2019, 2022b) methods.
(a) An example: patient 1 has disease $d_{3}$, and Patient 2 has disease $d_{4}$, their health states differ only in one disease, yet the prescribed medications differ significantly.
(b) The relationship between the similarity of the patient’s health state and the similarity of the medications taken by the patients.
Figure 1. The problem of insufficient personalization of medication recommendation.
[…]
The term ”health state” refers to the sum of all diseases and procedures within a single clinical visit. Figure 1 illustrates two real-life cases, explaining why personalized medication recommendation is crucial. In our study, we observed a common phenomenon, as depicted in Figure 1(a), where even if two patients may have highly similar health states, the required medications could be vastly different. Furthermore, we conducted pairwise comparisons of all medical records in MIMIC-III (Johnson et al., 2016), as shown in Figure 1(b), where the results indicate that among patients with medical record similarity ranging from 80% to 90%, the average similarity of prescribed medications was only 48.8%. In conclusion, due to factors such as individual patient differences, doctors often propose different treatment plans for similar diseases/procedures. Hence, personalized representations tailored to each patient are particularly important.
Previous research methods, such as (Yang et al., 2021a; Wu et al., 2022), emphasize the importance of prior medications, suggesting that current medication combinations be recommended based on medications prescribed during the last visit. Meanwhile, other methods, like (Yang et al., 2021b, 2023), advocate starting from a molecular perspective to construct more accurate representations of medications. However, these studies are primarily medication-centric and limited to addressing the aforementioned issue of personalization due to the following limitations:
(1) Failing to model a direct causal relationship between medications and patient’s specific health state: Most medications are specifically designed for one or two particular diseases/procedures, which means that the relationships between diseases/procedures and medications should be point-to-point causal relationships. Traditional methods typically match disease/procedure sets with medication sets, and construct set-to-set relationships based on co-occurrence. However, these co-occurrence-based methods generate excessive false correlations and fail to capture the direct causal relationship between medications and the specific patient health state, making it impossible to determine which specific disease each medication is treating.
(2) Overlooking the difference of diseases in various health states: The same disease may assume different roles in various patient health states: it may be the primary cause of other diseases during one visit and a secondary disease resulting from other diseases during another visit. This intricate pathological relationship among diseases implies that the impact of diseases on patients varies across different health states. A similar issue exists with procedures. In dealing with multiple diseases/procedures, existing methods simply sum them with equal weight, failing to consider the varying roles that diseases play in different health states.
To address the aforementioned limitations, based on causal inference (Hasan et al., 2023), we propose a patient health state-centric model named CausalMed. Specifically, first, to address limitation (1), we replace co-occurrence relationships with causal relationships. Through causal discovery, we eliminate backdoor paths (Gultchin et al., 2020; Li et al., 2022) between medical entities (diseases, procedures, medications). Subsequently, we construct corresponding graph networks by estimating the quantified therapeutic effects of specific medications on specific diseases. Ultimately, we transform the initially chaotic set-to-set relationships into clear point-to-point relationships, modeling the direct causal relationships between medications and specific patient states. At the same time, to address limitation (2), we introduce a novel approach named Dynamic Self-Adaptive Attention (DSA). Grounded in the causal-based pathological relationships of each health state, DSA learns the variability in the effects that medical entities have on patients under different health states, leading to the construction of personalized patient representations. Ultimately, by applying this method to historical medical visit data, we integrate longitudinal visit information, resulting in tailored medication recommendations.
Detailed case study can be found in subsection 5.6. Our source code is publicly available on GitHub [^1] https://github.com/lixiang-222/CausalMed. We list our main contributions as follows:
  • Important Discovery: We make the first discovery that different diseases express significantly different meanings under various patient health states, and current works fail to capture which specific disease each medication is treating. These factors contribute to the low level of personalization in existing medication recommendations.
  • Novel Framework: We propose a novel medication recommendation framework centered on patients’ clinical health state. Utilizing causal inference, we capture point-to-point relationships between diseases and medications and develop a Dynamic Self-Adaptive Attention (DSA) mechanism to discern the dynamic differences in diseases across various health states, thereby enhancing the personalization of patient representations.
  • Extensive Evaluation: Through extensive experiments on real-world datasets, we demonstrate our model’s ability to capture personalized patient representations, outperforming other state-of-the-art baselines significantly.
2. Related Work
2.1. Medication Recommendation
Early works in medication recommendation like LEAP (Zhang et al., 2017) are instance-based, prescribing on single-visit data, and viewing medication recommendation as a multi-instance and multi-label classification task.
The first category of research focuses on modeling the sequential relationships in patients’ longitudinal medical records to address issues of personalization. RETAIN (Choi et al., 2016b) models patients’ histories to personalize disease predictions. MICRON (Yang et al., 2021a) adjusts medications for new symptoms to optimize therapy and minimize side effects. COGNet (Wu et al., 2022) translates patient’s diseases/procedures to medications and reuses effective past treatments.
The second category of methods aims to capture the relationships between entities, thereby addressing the problem of low accuracy in prescriptions. DMNC (Le et al., 2018) combines dynamic memory and deep learning to analyze case data and recommend precise medications. Several studies (Zheng et al., 2021b, 2023; Wu et al., 2023) utilize graph networks to describe the connections between medical entities, enhancing the ability to capture relationships, through innovations in constructing graphs. StratMed(Li et al., 2023) stratifies different disease/procedure-medication relationships, enhancing the representation of sparse relationships to alleviate the long-tail problem in the data.
The third category of methods involves incorporating partial medication-related knowledge to address safety and reliability issues in medication combinations. SafeDrug (Yang et al., 2021b) uses molecular data to improve medication relationship characterization and reduce adverse DDIs. MoleRec (Yang et al., 2023) simulates molecular substructures for better DDI representation. Carmen (Chen et al., 2023) integrates patient history with molecular learning to differentiate similar molecules by function.
Our proposed model employs a patient health state-centric framework and enhances personalized patient representations through the incorporation of causal inference methods.
Figure 2. CausalMed framework: The upper section illustrates the process of learning the patient’s representation from a single clinical visit. The lower section represents the integration of information from multiple visits and the prediction of medication combinations.
[…]
2.2. Causal Inference in Recommender Systems
The incorporation of causal inference (Gao et al., 2024) into recommender systems(RS) still presents as a relatively novel concept, and ongoing research directions can be broadly categorized into three main categories. The primary objective of the first research category is to tackle data bias issues. For instance, methods such as the backdoor criterion employed by (Wei et al., 2021; Zheng et al., 2021a; Wang et al., 2022d) aim to eliminate confounding factors, addressing challenges like popularity bias and exposure bias in the dataset.
In the second category of work, exemplified by (Zhang et al., 2021; Xiao and Wang, 2022; Wang et al., 2022c), real-world data collection is coupled with inverse probability weighting and counterfactual methods. This involves introducing counterfactual interactions that contradict observations to fill data gaps or reduce noise.
Additionally, the third category of works focus on enhancing model interpretability and fairness. (Wang et al., 2022a; Si et al., 2022) introduce causal discovery and counterfactual approaches in RS to tackle the inherent opacity issues associated with deep learning models.
Our method utilizes causal inference to establish causal relationships with clear direction and precise quantification, effectively overcoming the inherent limitations of traditional methods that depend on co-occurrence relationships.
3. Problem Definition
3.1. Medical Entity
A medical entity refers to specific medical concepts that constitute medical data. We use $\mathcal{D}=\{d_{1},d_{2},\ldots\}$, $\mathcal{P}=\{p_{1},p_{2},\ldots\}$, and =$\mathcal{M}=\{m_{1},m_{2},\ldots\}$ to represent three sets of medical entities: diseases, procedures, and medications.
3.2. DDI Matrix
Our Drug-Drug Interaction (DDI) data is extracted from the Adverse Event Reporting Systems (Hoffman et al., 2014). We represent DDI information using a binary matrix ${\mathbf{M}^{ddi}}\in\{0,1\}^{|\mathcal{M}|\times|\mathcal{M}|}$, where $\mathbf{M}^{ddi}_{ij}=1$ indicates the presence of an interaction between medication $m_{i}$ and medication $m_{j}$. A high frequency of DDI suggests potential safety issues in recommended results.
3.3. Causal Discovery and Estimation
Causal discovery and estimation constitute two pivotal aspects of causal inference (Hasan et al., 2023). Causal discovery entails employing statistical tools and machine learning algorithms to unveil causal relationships between variables. Causal estimation underscores the quantification of causal effects between variables, relying on observed data.
3.4. Input and Output
The model ingests Electronic Health Records (EHR) of patients as its primary input. Each patient’s record, symbolized as $\mathcal{H}$, involves multiple visits =$\mathcal{H}=\{v_{1},v_{2},\ldots,v_{t}\}$. Each visit record $v_{t}$ encapsulates a tripartite dataset: diseases, procedures, and medications, expressed as $v_{t}=\{\mathcal{D}_{t},\mathcal{P}_{t},\mathcal{M}_{t}\}$, where $\mathcal{D}_{t}$, $\mathcal{P}_{t}$, and $\mathcal{M}_{t}$ are all multi-hot encoded using 0 and 1. The output of model, denoted as ^$\hat{\mathcal{M}}_{t}$, is a predicted medication combination for $v_{t}$.
4. The Proposed Model: CausalMed
As depicted in Figure 2, our model starts by extracting entity representations from a single clinical visit. In the relationship mining stage, we employ causal discovery and estimation to uncover pathological relationships among medical entities and quantify the point-to-point relationships. During the phase of health state learning, we capture the dynamic changes of entities in different patient states through causal relationships, updating and aggregating embeddings of all medical entities to generate visit representations. Finally, in the information integration stage, we predict medication combinations based on longitudinal patient visit records.
4.1. Entity Representation
In this stage, we encode entities within a single clinical visit. Specifically, we start by extracting the current disease $\mathcal{D}_{t}$, current procedure $\mathcal{P}_{t}$, and the previous medication $\mathcal{M}_{t-1}$. Following that, we construct embedding tables for diseases, procedures, and medications, represented as $\mathbf{E}_{d}\in\mathbb{R}^{|\mathcal{D}|\times dim}$, $\mathbf{E}_{p}\in\mathbb{R}^{|\mathcal{P}|\times dim}$, and $\mathbf{E}_{m}\in\mathbb{R}^{|\mathcal{M}|\times dim}$ respectively, where $dim$ represents the embedding dimension and each row in the table retains a vector for specific medical entity. Through the tables, we obtain embeddings for all entities.
(1) $\mathbf{h}_{d_{i}}=\mathbf{E}_{d}(d_{i}),\quad\mathbf{h}_{p_{j}}=\mathbf{E}_{p}(p_{j}),\quad\mathbf{h}_{m_{k}}=\mathbf{E}_{m}(m_{k}),$
where $d_{i}\subset\mathcal{D}_{t}$, $p_{j}\subset\mathcal{P}_{t}$, and $m_{k}\subset\mathcal{M}_{t-1}$ represent specific medical entities, $\mathbf{h}_{d_{i}}\in\mathbb{R}^{dim}$ represents the embedding for entity $d_{i}$, and $\mathbf{h}_{p_{j}}\in\mathbb{R}^{dim}$ and $\mathbf{h}_{m_{k}}\in\mathbb{R}^{dim}$ follow the same logic for entities $p_{j}$ and $m_{k}$, respectively.
4.2. Relationship Mining
Figure 3. An example of causal discovery, discovering the true relationship between $d_{1}$ and $d_{2}$ and eliminating the false relationship between $d_{2}$ and $m_{1}$, ultimately transforming set-to-set relationships into point-to-point relationships.
[…]
At this stage, we introduce methods of causal discovery and causal estimation to uncover the underlying genuine relationships among medical entities and eliminate spurious relationships caused by mere co-occurrence. We transform set-to-set relationships into precise point-to-point relationships and refine the causal effects for each group of relationships.
Firstly, as illustrated in Figure 3, the co-occurrence-based method simply associates all diseases and medications that appear together in a single medical visit, assuming that $d_{1}$ and $d_{2}$ are both related to $m_{1}$. However, causal discovery methods can learn that $d_{2}$ is caused by $d_{1}$ by resolving backdoor paths (Pearl and Mackenzie, 2018; Li et al., 2022), and there is no actual relationship between $m_{1}$ and $d_{2}$. We utilize causal discovery to learn the causal structure from set-to-set relationships, uncovering the true underlying relationships and eliminating spurious ones caused by co-occurrence, ultimately generating point-to-point relationships between diseases and medications.
Specifically, we analyze the EHR to obtain the distribution $U$ of medical entities and employ a variation of Greedy Intervention Equivalence Search (GIES) (Nandy et al., 2018; Kalainathan et al., 2020) to identify confounding bias and generate a causal relationship graph $G$. GIES optimizes the Bayesian equivalence class. Concerning the Bayesian equivalence class, we define the scoring criterion as $S(G,U)$ to evaluate the quality of each causal graph $G$ learned from the distribution $U$.
(2) $S(G,U)=\sum_{i=1}^{n}s(X_{i},P^{G}_{a_{i}}),$
(3) $G^{\prime}=\text{GIES}(S,G),$
where $n$ is the number of variables, $X_{i}$ is the variable including disease, surgery, and medication, and $P^{G}_{a_{i}}$ represents the parents of $X_{i}$ in the graph $G$. And $s(\cdot,\cdot)$ refers to the method of obtaining the equivalence class from $X_{i}$ and its parent nodes ($P^{G}_{a_{i}}$), through which we can produce the initial graph $G$ and continuously update it. The $\text{GIES}(\cdot)$ optimizes and learns an optimized graph $G^{\prime}$ from the initial graph $G$ and the equivalence score $S$. We apply the above process to each medical visit, generating a series of pathological graphs $\mathcal{G}=\{G_{v_{1}},G_{v_{2}},\ldots\}$ that encompass all visits.
Subsequently, adopting the principles of causal estimation, we define diseases and procedures as treatment variables and medications as outcome variables. We use a generalized linear model adjusted based on the backdoor criterion (Pearl and Mackenzie, 2018; Gultchin et al., 2020) to calculate the quantified effects of medications on diseases and procedures. Ultimately, we obtain a quantified relationship between each disease/procedure and each medication, and construct matrices of causal effects, denoted as $\mathbf{M}^{dm}\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{M}|}$ and $\mathbf{M}^{pm}\in\mathbb{R}^{|\mathcal{P}|\times|\mathcal{M}|}$.
4.3. Health State Learning
At this stage, we utilize causal relationships to learn the health state of patients. In homomorphic relationships, we capture the meanings of entities under different health states to update their representations; In heterogeneous relationships, we apply point-to-point disease/procedure-medication causality to update entity representations. Ultimately, we integrate this information to create a personalized representation of the clinical visit.
4.3.1. Homomorphic Relationship Learning
As shown in the left half of the Health State Learning part of Figure 2. In this stage, we primarily learn the representations of various entities through the relationships between homomorphic entities. By exploring pathological relationships based on causality, we delve into the dynamic differences between entities to enhance our ability to capture the personalized characteristics of patients.
In different health states, the impact of the same disease on patients can vary. In one state, a disease may be the cause of other diseases, while in another, it may be the result of them. Therefore, we introduce a Dynamic Self-Adaptive Attention (DSA) mechanism to learn the role of diseases in various health states. By using causal positioning to determine the role of a disease in the current clinical setting and assigning different weights, we enhance the personalized representation of patients.
First, we extract relationships from the causal graph $G_{v_{t}}$ that exclusively involves entities from the current visit and construct three isomorphic causal graphs: $G^{d}_{v_{t}}$, $G^{p}_{v_{t}}$, and $G^{m}_{v_{t-1}}$. And then we categorize the patient’s disease/procedure, as shown in Figure 4. taking the example of diseases, for the learning of differences between diseases in a clinical visit, we categorize the diseases $\mathcal{D}_{t}$ into four types $\mathcal{D}^{j}_{t}$ based on $G^{d}_{v_{t}}$: 1) Causal diseases $\mathcal{D}^{1}_{t}$, which initiate other diseases and represent the primary ailment; 2) Effect diseases $\mathcal{D}^{2}_{t}$, influenced by other diseases and indicative of secondary symptoms; 3) Middle diseases $\mathcal{D}^{3}_{t}$, both influencing and influenced by other diseases; 4) Independent diseases $\mathcal{D}^{4}_{t}$, existing without direct causal links to other diseases.
Since each patient has a unique disease history and record, the same disease may occupy different positions in the causal graphs $G^{d}_{v_{t}}$ of different patients, leading to classification into different disease groups. The DSA mechanism learns the dynamic changes of disease groups in different health states, effectively capturing the personalized differences of diseases when facing different patients. It updates the representations of each entity based on these groupings to enhance the expressiveness of key entities. The specific formula is as follows:
(4) $\mathcal{D}^{j}_{t}=\text{Classify}(d_{i},G^{d}_{v_{t}}),$
(5) $\mathbf{h}^{r}_{d_{i}}=\mathbf{h}_{d_{i}}\cdot w^{j}_{t},$
(6) $w^{j}_{t}=\frac{\exp(\mathbf{W}\cdot\mathbf{h}^{D^{j}}_{t}+b)}{\sum_{k=1}^{4}\exp(\mathbf{W}\cdot\mathbf{h}^{D^{k}}_{t}+b)},$
where $\text{Classify}(\cdot)$ assigns $d_{i}$ to a specific category $\mathcal{D}^{j}_{t}$, dynamic weight $w^{j}_{t}$ is generated by DSA, $\mathbf{W}$ and $b$ are trainable weight matrix and bias term, $\mathbf{h}^{D^{j}}_{t}$ represents the sum of embeddings of diseases within the ${D^{j}}_{t}$ set, and $\mathbf{h}^{r}_{d}$ is the modified embedding of diseases. We have extended the same approach to procedures and medications, obtained $\mathbf{h}^{r}_{p}$ and $\mathbf{h}^{r}_{m}$.
Figure 4. Using the classifier within DSA, we analyze the causal position of each node in the causal graph $G^{d}_{v_{t}}$ to gauge its influence on the patient’s current state. For instance, if $d_{3}$ is identified as a causal disease, it is classified as ${D}^{1}_{t}$.
[…]
4.3.2. Heterogeneous Relationship Learning
As shown in the right half of the Health State Learning part of Figure 2. At this stage, we update the representations of medical entities based on the correlations between diseases/procedures and medications obtained through causal estimation in relationship mining. This process aims to enhance the accuracy of entity representations.
We utilize the causal effect matrices $\mathbf{M}^{dm}$ and $\mathbf{M}^{pm}$ to construct disease-medication and procedure-medication complete bipartite graphs $G^{dm}_{v_{t}}$ and $G^{pm}_{v_{t}}$ for single visit.
Taking $G^{dm}_{v_{t}}$ as an example, the nodes encompass both $\mathcal{D}_{t}$ and $\mathcal{M}_{t-1}$. By utilizing Relational Graph Convolutional Networks (RGCN) (Schlichtkrull et al., 2018), we categorize edges corresponding to relationships with similar causal effects into the same type. Subsequently, we establish point-to-point relationships between diseases and medications. The representation of $d_{i}$ at layer $l$ is then updated on the point-to-point relationship using the following formula:
(7) $\mathbf{h}_{d_{i}}^{l+1}=\sigma(\mathbf{W}_{0}^{l}\cdot\mathbf{h}_{d_{i}}^{l}+\sum_{r\in\mathcal{R}}\mathbf{W}_{r}^{l}\cdot(\frac{1}{c_{{d_{i}},r}}\sum_{j\in\mathcal{N}_{r}({d_{i}})}\mathbf{h}_{j}^{l})),$
(8) $\mathbf{W}_{r}^{l}=I+\Theta_{r}^{l},$
where $\mathbf{W}_{r}^{l}$ is the weight matrix to relation $r$ at layer $l$, $\mathbf{W}_{0}^{l}$ is the weight matrix to node itself, $\mathcal{N}_{r}({d_{i}})$ is the set of neighboring nodes of ${d_{i}}$ with respect to relation $r$ and $c_{{d_{i}},r}$ is a normalization factor ensuring consistency in the influence of neighboring nodes under different relations. $I$ is the identity matrix, $\Theta_{r}^{l}$ is the weight update matrix for relation $r$ at layer $l$. In $G^{pm}_{v_{t}}$, we similarly repeat the aforementioned process, ultimately obtaining the entity embeddings $\mathbf{h}^{e}_{d}$, $\mathbf{h}^{e}_{p}$, and $\mathbf{h}^{e}_{m}$.
4.3.3. Aggregation of Entity Information
We obtain three set representations $\mathbf{h}_{\mathcal{D}_{t}}$, $\mathbf{h}_{\mathcal{P}_{t}}$, and $\mathbf{h}_{\mathcal{M}_{t-1}}$ by aggregating the updated embeddings with element-wise addition, which are then concatenated to the representation $\mathbf{h}_{v_{t}}$ for the current clinical visit.
(11) $\begin{gathered}\mathbf{h}_{\mathcal{D}_{t}}=\text{Agg}(\mathbf{h}^{r}_{d},\mathbf{h}^{e}_{d}),\mathbf{h}_{\mathcal{P}_{t}}=\text{Agg}(\mathbf{h}^{r}_{p},\mathbf{h}^{e}_{p}),\\ \mathbf{h}_{\mathcal{M}_{t-1}}=\text{Agg}(\mathbf{h}^{r}_{m},\mathbf{h}^{e}_{m}),\end{gathered}$
(12) $\mathbf{h}_{v_{t}}=[\mathbf{h}_{\mathcal{D}_{t}}||\mathbf{h}_{\mathcal{P}_{t}}||\mathbf{h}_{\mathcal{M}_{t-1}}].$
4.4. Medication Recommendation
To deep mine personalized patient representations, we consider the patient’s medical history. We represent each medical record and integrate this information to recommend medication combinations. Using multiple GRUs to capture temporal dependencies among visits, we generate the GRU output $\mathbf{o}_{v_{t}}$ and patient embedding $\mathbf{h}_{\mathcal{H}}$. An MLP layer with activation function $\sigma$ converts this representation into medication scores. Medication $\hat{m}_{i}$ with a score exceeding $\delta$ will be chosen.
(13) $\mathbf{h}_{\mathcal{H}}=\mathbf{o}_{v_{t}}=\text{GRU}(\mathbf{o}_{v_{t-1}},\mathbf{h}_{v_{t}}),$
(14) $score=\sigma(\text{MLP}(\mathbf{h}_{\mathcal{H}})),$
(15) $\hat{m}_{i}=\begin{cases}1,&score_{m_{i}}\geq\delta\\ 0,&score_{m_{i}}<\delta\end{cases}.$
Table 1. The performance of each model on the test set regarding accuracy and safety. The best and the runner-up results are highlighted in bold and underlined respectively under t-tests, at the level of 95% confidence level.
Model MIMIC-III MIMIC-IV
[…]
LR 0.4924 0.0830 0.6490 0.7548 16.0489 0.4569 0.0783 0.6064 0.6613 8.5746
ECC 0.4856 0.0817 0.6438 0.7590 16.2578 0.4327 0.0764 0.6129 0.6530 8.7934
RETAIN 0.4871 0.0879 0.6473 0.7600 19.4222 0.4234 0.0936 0.5785 0.6801 10.9576
LEAP 0.4526 0.0762 0.6147 0.6555 18.6240 0.4254 0.0688 0.5794 0.6059 11.3606
GAMENet 0.4994 0.0890 0.6560 0.7656 27.7703 0.4565 0.0898 0.6103 0.6829 18.5895
SafeDrug 0.5154 0.0655 0.6722 0.7627 19.4111 0.4487 0.0604 0.6014 0.6948 13.6943
MICRON 0.5219 0.0727 0.6761 0.7489 19.2505 0.4640 0.0691 0.6167 0.6919 12.7701
COGNet 0.5312 0.0839 0.6744 0.7708 27.6335 0.4775 0.0911 0.6233 0.6524 18.7235
MoleRec 0.5293 0.0726 0.6834 0.7746 22.0125 0.4744 0.0722 0.6262 0.7124 13.4806
CausalMed 0.5389 0.0709 0.6916 0.7826 20.5419 0.4899 0.0677 0.6412 0.7338 14.4295
4.5. Model Training and Inference
Our training and inference processes share the same pipeline. In training, we optimize all learnable parameters and define the following loss functions: binary cross-entropy loss $\mathcal{L}_{bce}$, multi-label margin loss $\mathcal{L}_{multi}$, and DDI loss $\mathcal{L}_{ddi}$.
(16) $\mathcal{L}_{bce}=-\sum_{i=1}^{|\mathcal{M}|}{m_{i}}\log({\hat{m}_{i}})+(1-{m_{i}})\log(1-{\hat{m}_{i}}),$
(17) $\mathcal{L}_{multi}=\sum_{i,j:{m_{i}}=1,{m_{j}}=0}\frac{\max(0,1-({\hat{m}_{i}}-{\hat{m}_{j}}))}{|\mathcal{M}|},$
(18) $\mathcal{L}_{ddi}=\sum_{i=1}^{|\mathcal{M}|}\sum_{j=1}^{|\mathcal{M}|}\mathbf{M}^{ddi}_{ij}\cdot{\hat{m}_{i}}\cdot{\hat{m}_{j}}.$
We adopt a loss aggregation approach that aligns with the methodology used in prior work (Yang et al., 2023).
(19) $\mathcal{L}=\alpha(\beta\mathcal{L}_{bce}+(1-\beta)\mathcal{L}_{multi})+(1-\alpha)\mathcal{L}_{ddi},$
(20) $\alpha=\begin{cases}1,&\text{DDI rate}\leq\gamma\\ \max\{0,1-\frac{\text{DDI rate}-\gamma}{kp}\},&\text{DDI rate}>\gamma\end{cases},$
where $\beta$ is hyperparameters, and the controllable factor $\alpha$ is relative to DDI rate, $\gamma\in(0,1)$ is a DDI acceptance rate and $kp$ is a correcting factor for the proportional signal.
5. Experiments
In this section, we conduct extensive experiments aimed at answering the following key research questions (RQ):
  • RQ1: Does CausalMed provide more accurate and safe recommendations than state-of-the-art medication recommendation systems?
  • RQ2: Can the two core modules proposed in this paper enhance the effectiveness of recommendations?
  • RQ3: Are the causal relationships used by CausalMed more appropriate for medication recommendations than the co-occurrence relationships used in previous medication recommendations?
  • RQ4: Why CausalMed can generate more personalized patient representations based on causality?
  • RQ5: Can the representation process of CausalMed be made transparent, and are its results interpretable?
5.1. Experiments Setup
Table 2. Statistics of the datasets.
Item MIMIC-III MIMIC-IV
# patients 6,350 60,125
# clinical events 15,032 156,810
# diseases 1,958 2,000
# procedures 1,430 1,500
# medications 131 131
avg. # of visits 2.37 2.61
avg. # of medications 11.44 6.66
5.1.1. Datasets
We verify the model performance on the MIMIC-III (Johnson et al., 2016) and MIMIC-IV (Johnson et al., 2023). Our methodology for data processing and evaluation mirrored that of (Yang et al., 2021b), encompassing: 1) Data preprocessing, with outcomes detailed in Table 2; 2) Division of datasets into training, validation, and test sets in a 4/6-1/6-1/6 ratio; 3) A sampling approach involving the extraction of test data from the test set via ten rounds of bootstrap sampling.
5.1.2. Evaluation Metrics
1)Jaccard: measures the overlap between predicted and actual sets, indicating similarity; 2) DDI-rate: Assesses safety by indicating the potential for harmful medication interactions. A lower value signifies a safer medication combination; 3)F1-Score: Balances precision and recall, a generic evaluation metric for recommendation tasks; 4)Precision-Recall AUC (PRAUC): Represents the model’s ability to distinguish between classes across various thresholds. 5)Avg.#Med: Reflects the average number of medications recommended per clinical visit and is used for reference only, not indicative of the recommendation system’s overall quality.
Table 3. The performance of multiple variants of CausalMed on evaluation metrics. The best and the runner-up results are highlighted in bold and underlined respectively under t-tests, at the level of 95% confidence level.
Model MIMIC-III MIMIC-IV
[…]
CausalMed w/o T 0.5369 0.0734 0.6900 0.7824 0.4878 0.0702 0.6405 0.7330
CausalMed w/o P 0.5324 0.0731 0.6860 0.7745 0.4838 0.0695 0.6355 0.7258
CausalMed w/o T+P 0.5339 0.0740 0.6873 0.7788 0.4847 0.0711 0.6373 0.7309
CausalMed 0.5389 0.0709 0.6916 0.7826 0.4899 0.0677 0.6412 0.7338
5.1.3. Baselines
LR (Indra et al., 2016) is a classifier for medication recommendation but doesn’t consider historical impacts on predictions. ECC (Read et al., 2011) improves multi-label classification by splitting it into multiple binary problems. RETAIN (Choi et al., 2016b) uses a reverse time attention mechanism to highlight influential historical events in sequential medical data. LEAP (Zhang et al., 2017) uses reinforcement learning to optimize medication strategies by predicting label dependencies and DDI. GAMENet (Shang et al., 2019) employs a graph-enhanced memory network to enhance the safety and efficacy of medication combinations. SafeDrug (Yang et al., 2021b) predicts DDI by analyzing structural similarities between medication molecules. MICRON (Yang et al., 2021a) uses a residual network to personalize medication recommendations by updating the latest combinations. COGNet (Wu et al., 2022) generates medication sets using cloning and prediction mechanisms, reintroducing significant past medications. MoleRec (Yang et al., 2023) analyzes molecular substructures to better predict medication properties and interactions.
5.1.4. Implementation Details
In our study, as shown in subsection 5.4, we configure embedding dimensions at $\text{dim}=64$. The RGCN architecture consists of a 2-layer graph neural network segmented into 5 groups. For GRU and MLP, a single-layer network with a 64-dimensional hidden layer and ReLU activation is employed. The parameters, consistent with previous work (Yang et al., 2023), include a dropout rate of 0.5, a threshold $\delta=0.5$, loss function parameters $\beta=0.95$, $kp=0.05$, and an acceptance rate $\gamma=0.06$. Training uses the Adam Optimizer with a learning rate of $lr=0.0005$ and $L2$ regularization with a coefficient of 0.005. Experiments are conducted on an Ubuntu machine with 30GB memory, 12 CPUs, and a 24GB NVIDIA RTX3090 GPU.
5.2. Performance Comparison (RQ1)
To demonstrate the effectiveness of CausalMed, we conducted comparative experiments with several state-of-the-art baselines, and the results are shown in Table 1. In the realm of traditional machine learning-based methods, both LR and ECC demonstrate sub-optimal performance in terms of accuracy and safety. Despite employing deep learning strategies, LEAP fails to consider patients’ longitudinal historical visit information, resulting in no significant improvement in outcomes. Although RETAIN and GAMENet effectively improve accuracy, they ignore the consideration of DDI issues. SafeDrug introduces the importance of molecules and uses molecular characterization to identify potential DDI problems, greatly improving safety. However, it fails to balance accuracy and safety, performing poorly in terms of accuracy. MICRON proposes using the residual network to learn the difference in patient status between two consecutive visits. It does not directly prescribe medication but updates the last prescription. The results are greatly improved. COGNet treats this task as a translation task and uses Transformer to further improve efficiency. However, due to insufficient consideration of safety, there are still certain flaws. MoleRec conducts further research on molecules, using the substructure of molecules to learn medication representations and recommendations, improving in multiple dimensions.
CausalMed introduces a health state-centric framework, recommending personalized medication combinations based on causality. Experimental results demonstrate a significant enhancement of our proposed method compared to other baselines, validating the effectiveness of CausalMed.
Table 4. The impact of three hyperparameters on the model. The best and the runner-up results are highlighted in bold and underlined respectively under t-tests, at the level of 95% confidence level. Parameters used in the final model are in bold.
parameter MIMIC-III MIMIC-IV
Edge type 4 5 6 7 4 5 6 7
[…]
RGCN layer 1 2 3 4 1 2 3 4
[…]
Embedding dim 32 64 128 256 32 64 128 256
[…]
5.3. Ablation Study (RQ2&RQ3&RQ4)
To verify the effectiveness of our proposed innovations, we thoroughly evaluate three variant models. CausalMed $w/o$ T omits the causal effects and the refinement of the point-to-point relationships between diseases/procedures and medications. CausalMed $w/o$ P excludes pathological relationships, thereby not identifying disease differences. CausalMed $w/o$ T+P represents the model with the concurrent removal of both modules. Since the ablation versions originate from the same framework, we remove the Avg.#Med which only served as a reference.
As illustrated in Table 3, the result indicates that compared to CausalMed $w/o$ T+P, both CausalMed $w/o$ T and CausalMed $w/o$ P show significant improvements in accuracy. This indicates that CausalMed $w/o$ T can identify causal point-to-point associations between diseases/surgeries and medications, providing a more accurate relational basis through the graph network process. At the same time, CausalMed $w/o$ P can learn personalized representations based on the health state of the patient. Notably, CausalMed $w/o$ P exhibits a significantly higher DDI rate. This is attributed to our modeling of inter-medication relationships during the health state learning stage, where we learn about the causal interactions between medications, thereby enhancing the safety of the outcomes. The higher performance of CausalMed indicates that the two modules have a synergistic effect when used together, leading to a more pronounced enhancement. Through the construction of these variant models, we can effectively verify the contributions of the different innovation points to our overall model.
5.4. Parameters Analysis
In this subsection, we investigate the impact of several hyperparameters. These parameters include the number of types of edges in the bipartite graph, the number of layers in the RGCN, and the embedding dimensions for each medical entity. We select four neighboring configurations and utilize the Jaccard and DDI rate to represent the accuracy and safety of the results in Table 4.
Our analysis indicates that optimal and balanced outcomes are achieved with a configuration of 5 edge types, 2 RGCN layers, and 64 dimensions. The number of edge types delineates the granularity of relationship categorization, where a limited number can result in too broad divisions, reducing the uniqueness of embeddings, and excess may lead to smaller sample sizes per category, compromising the embeddings’ representativeness. The two-layer RGCN facilitates the integration of the two node types within the graph, while additional layers could obscure the demarcation among embeddings. In terms of dimension, low dimensions risk insufficient training, and excessively high dimensions could impinge on both accuracy and safety, owing to overfitting.
5.5. Quality Study (RQ3)
(a) Results of the co-occurrence relationship-based model on the training and validation sets.
(b) Results of causality-based models on training and validation sets.
Figure 5. The performance of co-occurrence-based and causality-based methods on Jaccard.
[…]
To validate the significance of introducing causality, we compare the performance of causal relationships and co-occurrence relationships in our model. We create a comparative model by replacing all causal relationships in CausalMed with co-occurrence relationships. Specifically, we remove causal inference methods and replace the causal effects between diseases/procedures and medications with co-occurrence rates. Since co-occurrence methods cannot capture the pathological relationships between medical entities, we remove an entire process of health state learning and fuse them using equal-weight addition. The subsequent recommendation methods remain the same as the original model.
The results for both methods on training and validation sets are illustrated in Figure 5, indicating that the co-occurrence-based method is inferior to the causality-based method, whether on the training or validation set. Firstly, the causal effects between diseases and medications modeled through causal estimation are characterized by point-to-point, offering finer details than co-occurrence rates derived from set-to-set relationships. Furthermore, methods based on causal relationships can dynamically construct patient representation, whereas methods based on co-occurrence simply aggregate multiple entity embeddings. Therefore, methods based on causality are more suitable for medication recommendation compared to those based on co-occurrence relationships.
5.6. Case Study (RQ3&RQ4&RQ5)
(a) Quantitative relationships between specific diseases/adverse reactions and medications based on co-occurrence relationships.
(b) Quantitative relationships between specific diseases/adverse reactions and medications based on causal relationships.
(c) Prescription principles for a specific visit.
Figure 6. The learning process under co-occurrence-based and causality-based approaches from the real clinical sample.
[…]
To more intuitively demonstrate how causality-based methods capture personalized patient representations, we employ the same co-occurrence-based comparative model as in the previous subsection and make transparent the learning process of patient representations in a real case under both approaches. This ultimately proves the rationality and interpretability of the recommended results in this study. Figure 6 is derived from a clinical visit associated with gastric bleeding. Sub-figure 6(a) and 6(b) illustrate the quantified relationships between diseases/procedures and medications, while Sub-figure 6(c) provides a visualization of the learning results.
Sub-figure 6(a) shows co-occurrence-based relationships between diseases/procedures and medications, which are ambiguous. Medications like $m_{5}$ (analgesics), $m_{6}$ (intestinal anti-inflammatory agents), and $m_{7}$ (sulfonamides) are frequently prescribed and show high co-occurrence rates with many diseases. In contrast, rare medications like $m_{3}$ (mineral supplements) have low co-occurrence rates, making specific associations difficult to determine. As depicted in the upper half of Sub-figure 6(c), this results in common medications being associated with nearly every disease/procedure while rare medications fail to establish appropriate relationships. Consequently, the model preferentially recommends common medications, overlooking rare ones, sometimes leading to positive feedback during training and amplifying the error.
In contrast, relationships based on causality are clearer, as shown in Sub-figure 6(b). Point-to-point causal estimation links each medication to its specific diseases or procedures, avoiding spurious correlations found in co-occurrence-based relationships. The lower half of Sub-figure 6(c) shows that by analyzing differences in patients’ diseases/procedures, we create dynamic patient profiles. By integrating precise disease/procedure-medication relationships, we build more personalized patient representations. Our method learns that doctors only need to prescribe for $d_{4}$ (Regional enteritis) and $p_{1}$ (Small bowel endoscopy) related to $d_{1}$ (Gastrointest hemorr), without prescribing specifically for $d_{1}$. In contrast, the co-occurrence-based method prescribes all these medications for $d_{1}$.
We demonstrate how causality-based methods enhance medication recommendation personalization and validate our proposed framework’s reasonability, transparency and interpretability via comparative analysis.
6. Conclusion
We introduce a causality-based, patient health state-centric framework named CausalMed, which effectively addresses the issues of ambiguous relationships between diseases/procedures and medications and the challenge of capturing differences among diseases/ procedures, thereby significantly enhancing the personalization of patient representations. Extensive experiments on real-world datasets validate the efficacy of the proposed framework.
Acknowledgements.
This work was supported by the Innovation Capability Improvement Plan Project of Hebei Province (No. 22567637H), the S&T Program of Hebei(No. 236Z0302G), and HeBei Natural Science Foundation under Grant (No.G2021203010 & No.F2021203038).